### v4

@@ 3 $\bm{r}$-Coverage
For each $q\in Q$, let $D_{q}$ be the radius-$r$ disk centered at $q$. Note that a radius-$r$ disk contains $Q$ iff its center lies in the common intersection $\bigcap_{q\in Q}D_{q}$. Therefore, $\mathcal{CR}_{r}(Q)$ is the union of all radius-$r$ disks with centers in $\bigcap_{q\in Q}D_{q}$. Equivalently, $\mathcal{CR}_{r}(Q)$ is the Minkowski sum of $\bigcap_{q\in Q}D_{q}$ and the radius-$r$ disk centered at the origin. Note that $\bigcap_{q\in Q}D_{q}$ is convex. As the Minkowski sum of two convex bodies is also convex, we obtain that $\mathcal{CR}_{r}(Q)$ must be convex. ∎
4 Algorithm of Theorem 1
In this section, we prove Theorem 1. As discussed in Section 1, comparing to most of the previous work on the planar two-center problem, our algorithm is relatively simple, at least conceptually. However, proving its correctness is highly intricate and we devote the entire Section 5 to it.
Let $S$ be a set of $n$ points in the plane. Without loss of generality, we assume that the minimum enclosing disk of $S$ is the unit disk with equation $x^{2}+y^{2}\leq 1$. Choose a sufficiently large constant integer $c$ (say $c=100$). Set $\mathbb{Z}_{/c}=\{z/c:z\in\mathbb{Z}\}$ and $\theta=\frac{2\pi}{c}$. Define $\varGamma=\{(\cos i\theta,\sin i\theta):i\in[c]\}$, which is a set of $c$ unit vectors that evenly partition the perigon around the origin.
@@ 4 Algorithm of Theorem 1
For each point $p\in A$, we compute an optimal $p$-anchored solution for $S$ by Lemma 4. Among all these solutions, we take the one with smallest radius, and denote it by $\{D_{0},D_{0}^{\prime}\}$. Next, for each unit vector $\vec{\gamma}\in\varGamma$, we sort the points in $S$ along the direction $\vec{\gamma}$; let $S_{\vec{\gamma}}$ denote the corresponding sorted sequence. This completes the preprocessing of our algorithm.
@@
Given a number $r>0$, our goal is to determine whether there exist two radius-$r$ disks whose union covers $S$, and if yes, return such two disks. For a sequence $L$ of points in the plane, we define the maximal $r$-coverable prefix of $L$, denoted by $\varPhi_{r}(L)$, as the longest prefix of $L$ that is $r$-coverable. For a region $R$ in the plane, we let $L\cap R$ denote the subsequence (not necessarily contiguous) of $L$ consisting of the points inside $R$.
Algorithm 1 gives the entire algorithm. We begin with checking whether $r\geq\mathsf{rad}(D_{0})$ (note that $\mathsf{rad}(D_{0})=\mathsf{rad}(D_{0}^{\prime})$). If yes, we simply return $\{D_{0},D_{0}^{\prime}\}$ (Line 1). Otherwise, we consider the vectors in $\varGamma$ one by one (the for-loop in Lines 2-7). For each $\vec{\gamma}\in\varGamma$, we do the following. First, we compute $X$, the maximal $r$-coverable prefix of $S_{\vec{\gamma}}$, and construct its $r$-coverage $\mathcal{CR}_{r}(X)$. Then, we compute $Y$, the maximal $r$-coverable prefix of $S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(X)$, and construct $\mathcal{CR}_{r}(Y)$. Finally, we compute $Z=S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(Y)$. Now we partition $S$ into two subsets, $Z$ and $S\setminus Z$ (here we view $Z$ as a set instead of a sequence). We simply check whether $Z$ and $S\setminus Z$ are both $r$-coverable. If so, we find a radius-$r$ disk $E$ (resp., $E^{\prime}$) that covers $Z$ (resp., $S\setminus Z$) and return $\{E,E^{\prime}\}$ as the solution. If no solution is found after all vectors in $\varGamma$ are considered, then we return NO (Line 8).
@@
1:if $r\geq\mathsf{rad}(D_{0})$ then return $\{D_{0},D_{0}^{\prime}\}$
2:for every $\vec{\gamma}\in\varGamma$ do
3:     $X\leftarrow\varPhi_{r}(S_{\vec{\gamma}})$
4:     $Y\leftarrow\varPhi_{r}(S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(X))$
5:     $Z\leftarrow S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(Y)$
6:     if there exist two radius-$r$ disks $E\supseteq Z$ and $E^{\prime}\supseteq S\setminus Z$ then
7:         return $\{E,E^{\prime}\}$
8:return NO
Time analysis.
We first show that the preprocessing can be implemented in $O(n\log n)$ time. Finding a farthest pair $(a,b)$ of $S$ takes $O(n\log n)$ time. For each $p\in A$, computing the optimal $p$-anchored solution can be done in $O(n\log n)$ time by Lemma 4. For each $\vec{\gamma}\in\varGamma$, computing the sorted sequence $S_{\vec{\gamma}}$ takes $O(n\log n)$ time. Since $A$ and $\varGamma$ are both of constant size, the overall preprocessing time is $O(n\log n)$. Then we consider the time cost of the decision procedure.
Observation 11.
Line 3-5 of Algorithm 1 can be implemented in $O(n)$ time.
Proof.
For Line 3, computing $\varPhi_{r}(S_{\vec{\gamma}})$ can be done in linear time by Lemma 3 since points of $S_{\vec{\gamma}}$ are already sorted along the direction $\vec{\gamma}$. Specifically, we begin with $Q=\emptyset$ and insert points of $S_{\vec{\gamma}}$ to $Q$ one by one. We can maintain the circular hull $\alpha_{r}(Q)$ in $O(1)$ amortized time using the data structure of Lemma 3, as the insertions to $Q$ are monotone. Note that $Q$ is $r$-coverable iff $\alpha_{r}(Q)$ exists. We keep inserting points to $Q$ until $\alpha_{r}(Q)$ does not exist. At this point, we obtain the maximal $r$-coverable prefix of $S_{\vec{\gamma}}$.
For Line 4, we first compute the $r$-coverage $\mathcal{CR}_{r}(X)$. To this end, we compute the circular hull $\alpha_{r}(X)$, which can be done in linear time using Lemma 3, as discussed above. After that, $\mathcal{CR}_{r}(X)$ can be obtained in linear time by Corollary 9. Next, we need to compute $S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(X)$. To this end, since the $r$-coverage $\mathcal{CR}_{r}(X)$ is convex by Lemma 10 and $S_{\vec{\gamma}}$ is sorted along $\vec{\gamma}$, computing $S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(X)$ can be easily done in linear time, e.g., by sweeping a line perpendicular to $\vec{\gamma}$. With $S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(X)$ available, computing $\varPhi_{r}(S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(X))$ can again be done in linear time by Lemma 3.
Similarly to the above for Line 4, Line 5 can also be implemented in linear time. ∎
The above observation implies that each iteration of the for-loop in Algorithm 1 takes $O(n)$ time, since Lines 6-7 can be done in $O(n)$ time using the linear-time algorithm for the planar one-center problem [7, 12, 23, 24] (or alternatively, using Lemma 3). As $|\varGamma|=O(1)$, the for-loop has $O(1)$ iterations. As such, the total runtime of Algorithm 1 is $O(n)$.
@@ 5 Correctness of the algorithm
In this section, we show the correctness of Algorithm 1, which is the last missing piece for proving Theorem 1. Before giving the analysis in Section 5.2, we first prove two geometric lemmas in Section 5.1 that will be needed in our analysis.
@@ 5.1 Geometric lemmas
Let $S$ be a set of points in the plane. Recall in Section 2 the definition of a tight $r$-solution for $S$. The following lemma describes a property of a tight $r$-solution.
Figure 4: Illustration of the proof of Lemma 12.
Lemma 12.
Let $\{D,D^{\prime}\}$ be a tight $r$-solution for $S$, and $\ell$ (resp., $\sigma$) be the line (resp., open segment) connecting the two disk centers $\mathsf{ctr}(D)$ and $\mathsf{ctr}(D^{\prime})$. Then, there exist two points $u,v\in(S\cap\partial D)\backslash D^{\prime}$ such that $\𝜎\overline{uv}\cap(\ell\backslash\sigma)\neq\emptyset$; see Figure 4. Similarly, there exist two points $u^{\prime},v^{\prime}\in(S\cap\partial D^{\prime})\backslash D$ such that $\𝜎\overline{u^{\prime}v^{\prime}}\cap(\ell\backslash\sigma)\neq\emptyset$.
Proof.
@@
Let $\ell^{\prime}$ be the line perpendicular to $\ell$ and passing through $\mathsf{ctr}(D)$. See Figure 4. We define $u$ as a point in $(S\cap\partial D)\backslash D^{\prime}$ on the side of $\ell^{\prime}$ not containing $\sigma$, i.e., the left side of $\ell^{\prime}$ by our assumption. Such a point always exists since $\{D,D^{\prime}\}$ is a tight $r$-solution of $S$; otherwise, we could move $D$ towards $D^{\prime}$ so that $\{D,D^{\prime}\}$ is still an $r$-solution while the distance between $\mathsf{ctr}(D)$ and $\mathsf{ctr}(D^{\prime})$ decreases, a contradiction. If there are multiple such points, we choose a leftmost one. Next, we choose $v$ as follows. If $u$ lies on $\ell$, we set $v=u$. In this case, $\overline{uv}$ intersects $\ell\setminus\sigma$ at the point $v$, and thus the lemma trivially holds. Below we assume $u\not\in\ell$.
@@
The above finds two different points $u$ and $v$ in $(S\cap\partial D)\backslash D^{\prime}$. Now we prove $\𝜎\overline{uv}\cap(\ell\backslash\sigma)\neq\emptyset$. Note that $\overline{uv}\cap\sigma=\emptyset$, due to our choice of $v$ (recall that $v$ is on $\partial D$ that lies on the side of $\overline{u\hat{u}}$ not containing $\sigma$ and $v\not\in\overline{u\hat{u}}$). Hence, it suffices to prove $\overline{uv}\cap\ell\neq\emptyset$. Assume to the contrary that $\overline{uv}\cap\ell=\emptyset$. Then, $u$ and $v$ lie on the same side of $\ell$. As $v$ is a point on $\partial D$ that lies on the side of $\overline{u\hat{u}}$ not containing $\sigma$, $v$ must be strictly to the left of $u$. But this contradicts our choice of $u$, i.e., $u$ is a leftmost point among all points of $(S\cap\partial D)\backslash D^{\prime}$ to the left of $\ell^{\prime}$. This proves $\overline{uv}\cap\ell\neq\emptyset$. ∎
@@
Let $\hat{u}$ and $\hat{v}$ be the antipodal points of $u$ and $v$ on $\partial D$, respectively. Also, let $q$ be the (unique) point on $\partial D$ closest to $\mathsf{ctr}(D^{\prime})$. Denote by $\tau$ the arc between $\hat{u}$ and $\hat{v}$ on $\partial D$ that does not contain $u$ and $v$. Since $\𝜎\overline{uv}\cap(\ell\backslash\sigma)\neq\emptyset$, such an arc $\tau$ must exist and $q\in\tau$. See Figure 5 for an illustration.
Figure 5: Illustration of the proof of Lemma 13.
Let $\tau^{\prime}$ be the antipodal arc of $\tau$ on $\partial D$, i.e., $\tau^{\prime}$ consists of the antipodal points on $\partial D$ of all points of $\tau$. Note that $u$ and $v$ are two endpoints of $\tau^{\prime}$. Since $\𝜎\overline{uv}\cap(\ell\backslash\sigma)\neq\emptyset$, $\tau^{\prime}$ is in a half circle of $\partial D$, and thus $\tau^{\prime}$ must be an arc of the circular hull $\alpha_{r}(\{u,v\})$ of $\{u,v\}$. Applying Corollary 7 with $Q=\{u,v\}$, we obtain that $\tau\subseteq\partial\mathcal{CR}_{r}(\{u,v\})$.
We claim that $(\partial D)\cap D^{\prime}\subseteq\tau$. Indeed, the four points $u,v,\hat{u},\hat{v}$ partition $\partial D$ into four arcs, one of which is $\tau$. Note that $u,v,\hat{u},\hat{v}\notin D^{\prime}$ by the assumption in the lemma statement. Therefore, $(\partial D)\cap D^{\prime}$ lies in one of these four arcs. We have $q\in(\partial D)\cap D^{\prime}$, provided that $(\partial D)\cap D^{\prime}\neq\emptyset$. As $q\in\tau$, $(\partial D)\cap D^{\prime}\subseteq\tau$ holds.
The above implies $(\partial D)\cap D^{\prime}\subseteq\tau\subseteq\partial\mathcal{CR}_{r}(\{u,v\})$, and hence $(\partial D)\cap D^{\prime}\subseteq(\partial\mathcal{CR}_{r}(\{u,v\}))\cap D^{\prime}$. On the other hand, we have $D\cap(\partial D^{\prime})\subseteq\mathcal{CR}_{r}(\{u,v\})\cap(\partial D^{\prime})$ since $\mathcal{CR}_{r}(\{u,v\})$ contains $D$. Note that for two closed convex bodies $B_{1}$ and $B_{2}$, it holds that $\partial(B_{1}\cap B_{2})=((\partial B_{1})\cap B_{2})\cup(B_{1}\cap(\partial B_{2}))$. Therefore, $\partial(D\cap D^{\prime})=((\partial D)\cap D^{\prime})\cup(D\cap(\partial D^{\prime}))$ and $\partial(\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime})=((\partial\mathcal{CR}_{r}(\{u,v\}))\cap D^{\prime})\cup(\mathcal{CR}_{r}(\{u,v\})\cap(\partial D^{\prime}))$, because $D,D^{\prime},\mathcal{CR}_{r}(\{u,v\})$ are all closed and convex (the convexity of $\mathcal{CR}_{r}(\{u,v\})$ follows from Lemma 10). It then follows that $\partial(D\cap D^{\prime})\subseteq\partial(\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime})$.
Now we are ready to prove $\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime}\subseteq D\cap D^{\prime}$. We distinguish two cases: whether $D\cap D^{\prime}$ is empty or not.
  • We first assume $D\cap D^{\prime}\neq\emptyset$. By Lemma 10, $\mathcal{CR}_{r}(\{u,v\})$ is convex. Thus, $\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime}$ is also convex. Since $D\cap D^{\prime}$ is convex and $\partial(D\cap D^{\prime})\subseteq\partial(\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime})$, it follows that $\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime}=D\cap D^{\prime}$ — it is not difficult to see that for any two (nonempty) convex bodies $B_{1}$ and $B_{2}$, if $\partial B_{1}\subseteq\partial B_{2}$, then $B_{1}=B_{2}$.
  • In the case $D\cap D^{\prime}=\emptyset$, we need to prove $\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime}=\emptyset$. Assume to the contrary that $\mathcal{CR}_{r}(\{u,v\})$ contains a point $p\in D^{\prime}$. Then, we have $D\cup\{p\}\subseteq\mathcal{CR}_{r}(\{u,v\})$. Note that, the convex hull of $D\cup\{p\}$ contains $q$ in its interior (by the definition of $q$). Furthermore, $\mathcal{CR}_{r}(\{u,v\})$, which is convex by Lemma 10, must contain the convex hull of $D\cup\{p\}$. We then have that $q$ must be in the interior of $\mathcal{CR}_{r}(\{u,v\})$. But this contradicts the fact that $q$ is a point on $\partial\mathcal{CR}_{r}(\{u,v\})$ since $q\in\tau$ and $\tau\subseteq\partial\mathcal{CR}_{r}(\{u,v\})$.
The lemma thus follows. ∎

### v5

@@ 3 $\boldsymbol{r}$-Coverage
For each $q\in Q$, let $D_{q}$ be the radius-$r$ disk centered at $q$. Note that a radius-$r$ disk contains $Q$ if and only if its center lies in the common intersection $\bigcap_{q\in Q}D_{q}$. Therefore, $\mathcal{CR}_{r}(Q)$ is the union of all radius-$r$ disks with centers in $\bigcap_{q\in Q}D_{q}$. Equivalently, $\mathcal{CR}_{r}(Q)$ is the Minkowski sum of $\bigcap_{q\in Q}D_{q}$ and the radius-$r$ disk centered at the origin. Note that $\bigcap_{q\in Q}D_{q}$ is convex. As the Minkowski sum of two convex bodies is also convex, we obtain that $\mathcal{CR}_{r}(Q)$ must be convex.
4 Algorithm of Theorem 1.1
In this section, we prove Theorem 1.1. As discussed in Section 1, compared to most of the previous work on the planar two-center problem, our algorithm is relatively simple, at least conceptually. However, proving its correctness is highly intricate and we devote the entire Section 5 to it.
Let $S$ be a set of $n$ points in the plane. Without loss of generality, we assume that the minimum enclosing disk of $S$ is the unit disk with equation $x^{2}+y^{2}\leq 1$. Choose a sufficiently large constant integer $c$ (say $c=100$). Set $\mathbb{Z}_{/c}=\{z/c:z\in\mathbb{Z}\}$ and $\theta=\frac{2\pi}{c}$. Define $\varGamma=\{(\cos i\theta,\sin i\theta):i\in[c]\}$, which is a set of $c$ unit vectors that evenly partition any disk centered at the origin.
@@
For each point $p\in A$, we compute an optimal $p$-anchored solution for $S$ by Lemma 2.2. Among all these solutions, we take the one with smallest radius, and denote it by $\{D_{0},{D_{0}}^{\prime}\}$. Next, for each unit vector $\vec{\gamma}\in\varGamma$, we sort the points in $S$ along the direction $\vec{\gamma}$; let $S_{\vec{\gamma}}$ denote the corresponding sorted sequence. This completes the preprocessing of our algorithm.
@@
Given a number $r>0$, our goal is to determine whether there exist two radius-$r$ disks whose union covers $S$, and if yes, return two such disks. For a sequence $L$ of points in the plane, we define the maximal $r$-coverable prefix of $L$, denoted by $\varPhi_{r}(L)$, as the longest prefix of $L$ that is $r$-coverable. For a region $R$ in the plane, we let $L\cap R$ denote the subsequence (not necessarily contiguous) of $L$ consisting of the points inside $R$.
Algorithm 1 gives the entire algorithm. We begin with checking whether $r\geq\mathsf{rad}(D_{0})$ (note that $\mathsf{rad}(D_{0})=\mathsf{rad}({D_{0}}^{\prime})$). If yes, we simply return $\{D_{0},{D_{0}}^{\prime}\}$ (Line 1). Otherwise, we consider the vectors in $\varGamma$ one by one (the for-loop in Lines 1-1). For each $\vec{\gamma}\in\varGamma$, we do the following. First, we compute $X$, the maximal $r$-coverable prefix of $S_{\vec{\gamma}}$, and construct its $r$-coverage $\mathcal{CR}_{r}(X)$. Then, we compute $Y$, the maximal $r$-coverable prefix of $S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(X)$, and construct $\mathcal{CR}_{r}(Y)$. Finally, we compute $Z=S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(Y)$. Observe that $X\subseteq Y\subseteq Z$. Now we partition $S$ into two subsets, $Z$ and $S\setminus Z$ (here we view $Z$ as a set instead of a sequence). We simply check whether $Z$ and $S\setminus Z$ are both $r$-coverable. If so, we find a radius-$r$ disk $E$ (resp., $E^{\prime}$) that covers $Z$ (resp., $S\setminus Z$) and return $\{E,E^{\prime}\}$ as the solution. If no solution is found after all vectors in $\varGamma$ are considered, then we return NO (Line 1). Figure 4 illustrates the algorithm with a simple example.
Time analysis.
We first show that the preprocessing can be implemented in $O(n\log n)$ time. Finding a farthest pair $(a,b)$ of $S$ takes $O(n\log n)$ time. For each $p\in A$, computing the optimal $p$-anchored solution can be done in $O(n\log n)$ time by Lemma 2.2. For each $\vec{\gamma}\in\varGamma$, computing the sorted sequence $S_{\vec{\gamma}}$ takes $O(n\log n)$ time. Since $A$ and $\varGamma$ are both of constant size, the overall preprocessing time is $O(n\log n)$. Then we consider the time cost of the decision procedure.
1 if $r\geq\mathsf{rad}(D_{0})$ then
2      return $\{D_{0},{D_{0}}^{\prime}\}$
3      for every $\vec{\gamma}\in\varGamma$ do
4           $X\leftarrow\varPhi_{r}(S_{\vec{\gamma}})$;
5           $Y\leftarrow\varPhi_{r}(S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(X))$;
6           $Z\leftarrow S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(Y)$;
7           if there exist two radius-$r$ disks $E\supseteq Z$ and $E^{\prime}\supseteq S\setminus Z$ then
8                return $\{E,E^{\prime}\}$
9
return NO
@@ 6           $Z\leftarrow S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(Y)$;
{cfigure}
[ht]
Illustration of Algorithm 1. (a) The red points are not in $\mathcal{CR}_{r}(X)$: $X$ consists of the three points inside the dashed circle. (b) The blue point is not in $\mathcal{CR}_{r}(Y)$: $Y$ consists of the four points inside the dashed circle. (c) The points of $Z$ and $S\setminus Z$ are covered by two radius-$r$ disks, respectively. Observe that $X\subseteq Y\subseteq Z$.
Lemma 4.1.
Lines 1-1 of Algorithm 1 can be implemented in $O(n)$ time.
Proof 4.2.
For Line 1, computing $\varPhi_{r}(S_{\vec{\gamma}})$ can be done in linear time by Lemma 2.1 since points of $S_{\vec{\gamma}}$ are already sorted along the direction $\vec{\gamma}$. Specifically, we begin with $Q=\emptyset$ and insert points of $S_{\vec{\gamma}}$ to $Q$ one by one. We can maintain the circular hull $\alpha_{r}(Q)$ in $O(1)$ amortized time using the data structure of Lemma 2.1, as the insertions to $Q$ are monotone. Note that $Q$ is $r$-coverable if and only if $\alpha_{r}(Q)$ exists. We keep inserting points to $Q$ until $\alpha_{r}(Q)$ does not exist. At this point, we obtain the maximal $r$-coverable prefix of $S_{\vec{\gamma}}$.
For Line 1, we first compute the $r$-coverage $\mathcal{CR}_{r}(X)$. Recall that we can compute $\alpha_{r}(X)$ in linear time using Lemma 2.1 as discussed above. After that, $\mathcal{CR}_{r}(X)$ can be obtained from $\alpha_{r}(X)$ in linear time by Corollary 3.7. Next, we need to compute $S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(X)$. To this end, since the $r$-coverage $\mathcal{CR}_{r}(X)$ is convex by Lemma 3.8 and $S_{\vec{\gamma}}$ is sorted along $\vec{\gamma}$, computing $S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(X)$ can be easily done in linear time, e.g., by sweeping a line perpendicular to $\vec{\gamma}$. With $S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(X)$ available, computing $\varPhi_{r}(S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(X))$ can again be done in linear time by Lemma 2.1.
Similarly to the above for Line 1, Line 1 can also be implemented in linear time.
The above observation implies that each iteration of the for-loop in Algorithm 1 takes $O(n)$ time, since Lines 1-1 can be done in $O(n)$ time using the linear-time algorithm for the planar one-center problem [ref:ChazelleOn96, ref:DyerOn86, ref:MegiddoLi83, Megiddo-linear], or alternatively, using Lemma 2.1. As $|\varGamma|=O(1)$, the for-loop has $O(1)$ iterations. As such, the total runtime of Algorithm 1 is $O(n)$.
@@ 5 Correctness of the algorithm
In this section, we show the correctness of Algorithm 1, which is the last missing piece for proving Theorem 1.1. Before giving the analysis in Section 5.2, we first prove two geometric lemmas in Section 5.1 that will be needed in our analysis.
@@ 5.1 Geometric lemmas
Let $S$ be a set of points in the plane. Recall the definition of a tight $r$-solution for $S$ in Section 2. The following lemma describes a property of a tight $r$-solution (similar observations were made previously, e.g., [Sharir-twocenter]).
Figure 4: Illustration of the proof of Lemma 5.1.
Lemma 5.1.
Let $\{D,D^{\prime}\}$ be a tight $r$-solution for $S$, and $\ell$ (resp., $\sigma$) be the line (resp., open segment) connecting the two disk centers $\mathsf{ctr}(D)$ and $\mathsf{ctr}(D^{\prime})$. Then, there exist two (possibly the same) points $u,v\in(S\cap\partial D)\backslash D^{\prime}$ such that $\𝜎\overline{uv}\cap(\ell\backslash\sigma)\neq\emptyset$; see Figure 4. Similarly, there exist two (possibly the same) points $u^{\prime},v^{\prime}\in(S\cap\partial D^{\prime})\backslash D$ such that $\𝜎\overline{u^{\prime}v^{\prime}}\cap(\ell\backslash\sigma)\neq\emptyset$.
Proof 5.2.
@@
Let $\ell^{\prime}$ be the line perpendicular to $\ell$ and passing through $\mathsf{ctr}(D)$. See Figure 4. We define $u$ as a point in $(S\cap\partial D)\backslash D^{\prime}$ on the side of $\ell^{\prime}$ not containing $\sigma$, i.e., the left side of $\ell^{\prime}$ by our assumption. Such a point always exists since $\{D,D^{\prime}\}$ is a tight $r$-solution of $S$; otherwise, we could move $D$ towards $D^{\prime}$ so that $\{D,D^{\prime}\}$ is still an $r$-solution while the distance between $\mathsf{ctr}(D)$ and $\mathsf{ctr}(D^{\prime})$ decreases, a contradiction. If there are multiple such points, we choose a leftmost one. Next, we choose $v$ as follows. If $u$ lies on $\ell$, we set $v=u$. In this case, $\overline{uv}$, which refers to a single point, intersects $\ell\setminus\sigma$ at the point $v$, and thus the lemma trivially holds. Below we assume $u\not\in\ell$.
@@
The above finds two different points $u$ and $v$ in $(S\cap\partial D)\backslash D^{\prime}$. Now we prove $\𝜎\overline{uv}\cap(\ell\backslash\sigma)\neq\emptyset$. Note that $\overline{uv}\cap\sigma=\emptyset$, due to our choice of $v$ (recall that $v$ is on $\partial D$ that lies on the side of $\overline{u\hat{u}}$ not containing $\sigma$ and $v\not\in\overline{u\hat{u}}$). Hence, it suffices to prove $\overline{uv}\cap\ell\neq\emptyset$. Assume to the contrary that $\overline{uv}\cap\ell=\emptyset$. Then, $u$ and $v$ lie on the same side of $\ell$. As $v$ is a point on $\partial D$ that lies on the side of $\overline{u\hat{u}}$ not containing $\sigma$, $v$ must be strictly to the left of $u$. But this contradicts our choice of $u$, i.e., $u$ is a leftmost point among all points of $(S\cap\partial D)\backslash D^{\prime}$ to the left of $\ell^{\prime}$. This proves $\overline{uv}\cap\ell\neq\emptyset$.
@@
Figure 5: Illustration of the proof of Lemma 5.3. The gray region is $D^{\prime\prime}\setminus D$. The two red regions are $\mathcal{CR}_{r}(\{u,v\})$ excluding $D\cup D^{\prime\prime}$. $\mathcal{CR}_{r}(\{u,v\})\setminus D$ consists of the gray region and the two red regions.
Let $\hat{u}$ and $\hat{v}$ be the antipodal points of $u$ and $v$ on $\partial D$, respectively. By Corollary 3.7, $\mathcal{CR}_{r}(\{u,v\})\setminus D$ is a subset of the union of the following parts: (1) a disk $D^{\prime\prime}$, not $D$, of radius $r$ with $u$ and $v$ on its boundary, and (2) two circular sectors centered at $u$ and $v$ of radius $2r$ and tangent to $D$ and $D^{\prime\prime}$ at their antipodal points of $u$ and $v$, respectively (in Figure 5, the red regions are the two circular sectors excluding the points in $D$). Let $\ell^{\prime}$ be the bisector of the two disk centers $\mathsf{ctr}(D)$ and $\mathsf{ctr}(D^{\prime})$. Note that $u,v,\hat{u},$ and $\hat{v}$ lie on $\partial D$ but not in $D^{\prime}$ by the assumption in the lemma statement. By definition, the union of the above parts (1) and (2) completely lies in the side of $\ell^{\prime}$ containing $\mathsf{ctr}(D)$. Therefore, $\mathcal{CR}_{r}(\{u,v\})\setminus D$ also completely lies in the side of $\ell^{\prime}$ containing $\mathsf{ctr}(D)$. Hence, no point of $\mathcal{CR}_{r}(\{u,v\})\setminus D$ lies in the side of $\ell^{\prime}$ containing $\mathsf{ctr}(D^{\prime})$. On the other hand, every point of $D^{\prime}$ in the side of $\ell^{\prime}$ containing $\mathsf{ctr}(D)$ must be in $D$. It thus follows that $\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime}\subseteq D\cap D^{\prime}$.